OpenStudy (anonymous):

Which number is between 0.5 and 5/8?

OpenStudy (anonymous):

What is your answer choice?

OpenStudy (anonymous):

1/4 8/28 15/28 19/28

OpenStudy (anonymous):

5/8=0.625

OpenStudy (anonymous):

0.5=1/2 =14/28 so its b

OpenStudy (anonymous):

do you agree?

OpenStudy (anonymous):

But, how is that in-between?

OpenStudy (anonymous):

That would be equal right?

OpenStudy (anonymous):

as the value of 15/28 which is 0.535714285 is between 0.5 and 5/8 which is 0.625 in decimal.

OpenStudy (anonymous):

It is 15/28= Which would be answer C not B

OpenStudy (anonymous):

14/28 is not even a answer choice

OpenStudy (anonymous):

I read wrong but at least I wrote it out

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

ok so im confused. Which is the answer?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

15/28

OpenStudy (anonymous):

15/28 is the only answer choice that is between the 2 numbers

OpenStudy (anonymous):

Thank you guys sooooo much!!! Could you do some more with me?

OpenStudy (anonymous):

If you tag me I will be there